OpenStudy (anonymous):

Express 6cosx-sinx in the form Rcos(x+a), where R>0 and 0

OpenStudy (campbell_st):

1st find the R value \[R = \sqrt{6^2 + (-1)^2} = \sqrt{37}\] then \[\sqrt{37} (6/\sqrt{37} \cos(x) - 1/\sqrt{37} \sin(x))\] then using the sum of 2 angle2 expansion for cos \[\cos(x + \alpha) = \cos(x)\cos(\alpha) - \sin(x) \sin (\alpha)\] then \[\cos(\alpha) = 6/\sqrt{37}\] \[\sin(\alpha) = 1/\sqrt{37}\] tan = sin/cos then \[\tan \alpha = 1/6\] \[\sqrt{6}\cos(x) - \sin(x) = \sqrt{37} \cos(x + 0.165)\] than angle is expressed in radians

OpenStudy (anonymous):

I understand everything except the first part. Why is it (-1)?

OpenStudy (campbell_st):

because the double angle expansion of cos(a + b) is cos(a)cos(b) - sin(a)sin(b) I hope that makes sense

OpenStudy (anonymous):

Ok, I know that's the expansion, but how did you arrive at (-1)^2?

OpenStudy (campbell_st):

its pythagoras's theorem R^2 = a^2 + b^2 a = 6 and b = -1

OpenStudy (anonymous):

Ah OK. :) Thanks

OpenStudy (anonymous):

Could you help with the latest one I put up?

OpenStudy (campbell_st):

I did the derivative...

OpenStudy (anonymous):

Of which?